The perimeter of a rectangle is 154 km and the length is 11 km more than the width. Find length and width.
Answer by addingup(3677)

2L+2W = 154
and
L = W+11 Substitute above:
2(W+11)+2W = 154
2W+22+2W = 154
4W = 132
W = 33 and L = W+11 = 33+11 = 44
-----------------------------
Check:
2(33)+2(44) = 154
66+88 = 154 Correct.
:
John